OpenStudy (anonymous):

If you know of an arithmetic sequence X1 + X3 + X5 + X7 + X9 + X11 = 144, then X1 + X6 + X11 = . . .

hartnn (hartnn):

72 its simple.

hartnn (hartnn):

6a+30d=144 so 3a+15d =72

hartnn (hartnn):

X1 + X6 + X11 = 3a+15d

OpenStudy (anonymous):

\[a+a+2d+a+4d+a+6d+a+8d+a+10d=144\]\[6a+30d=144\] \[3a+15d=\]